OpenStudy (anonymous):

Find the value of c that makes the expression a perfect square trinomial. Then write the expression as the square of a binomial. x^2+7x+c Please show work because i don't understand it.

OpenStudy (anonymous):

answer: 49/4 i think..

OpenStudy (anonymous):

Can you show work please??? I think that is the answer but idk how to get it.

OpenStudy (anonymous):

because the formula for a perfect square trinomial is a^2+2ab+(b/2)^2???

OpenStudy (anonymous):

if you work the answer backward (x+7/2)(x+7/2) =x^2+2(7/2)+(7/2)^2 =X^2+7+(49/4)

OpenStudy (anonymous):

there is a process called "completing the square" which will give the answer, btw 49/4 is the correct answer. but the process is this: in the expression\[x^2+bx+c\]the value of c for the poly to be a perfect square trinomial is\[\left( \frac{b}{2} \right)^2\]. in your example b=7, so c needs to be\[\left( \frac{7}{2} \right)^2=\frac{49}{4}\]

OpenStudy (anonymous):

btw cool green eye

OpenStudy (anonymous):

Thanks, help me with another!!! x^2-13x+c Same concept?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[c=\left( \frac{13}{2} \right)^2=\frac{169}{4}\]so the perfect square trinomial expression is\[x^2-13x+\frac{169}{4}\]

OpenStudy (anonymous):

add half of the middle coefficient, squared

OpenStudy (anonymous):

wow! the server just kicked me off, but i'm back now. as i was trying to say: this method only works if the first coefficient (the with x^2) is a +1

OpenStudy (anonymous):

okay, i think i got it :) I have a word problem, do u think you can help me?

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

Okay you have to use completing the square to find the vertex in the funtion's graph.

OpenStudy (anonymous):

ok shoot

OpenStudy (anonymous):

When you walk x meters per minute, your rate y of energy use (in calories per minute) can be modeled by y=0.0085x^2-1.5x+120.

OpenStudy (anonymous):

\[y=0.0085x^2-1.5x+120\] we need the leading coefficient to be a 1, and we need to get the 120 on the left side temporarily: first subtract 120, then divide by 0.0085\[\frac{y-120}{0.0085}=x^2-\left( \frac{1.5}{.0085} \right)x\]

OpenStudy (anonymous):

now we will multiply the coefficient for the x-term by 1/2 and then square it and add it the right side to make a perfect square trinomial.

OpenStudy (anonymous):

here's the catch: this is an equation and if we add this number to the right side, then we must also add it to the left side, right? OR, we can add and subtract it to the right side only. same difference, no?

OpenStudy (anonymous):

\[\frac{y-120}{.0085}=x^2-\left( \frac{1.5}{.0085} \right)x+\left( \frac{1}{2}*\frac{1.5}{.0085} \right)^2-\left( \frac{1}{2}*\frac{1.5}{.0085} \right)^2\]the first three terms on the right side is a perfect square trinomial; the subtracted term will (in part) be the y-coordinate of the vertex.

OpenStudy (anonymous):

Alright, we can move on ^_^ Yayy! :D

OpenStudy (anonymous):

i forgot to take the square root to get the 88.235...\[\frac{y-120}{.0085}\approx(x-88.2353)^2-1767785.467\]now all we have to do is solve for y

OpenStudy (anonymous):

i don't know if this happening to anyone else, but i lost my connection to the open study server again; i'm going to finish the problem in hope you'll come to view it :(

OpenStudy (anonymous):

the standard form of a quadratic function with vertex (h, k) is\[y=a(x-h)^2+k\]compare this to the equation from your problem and you can see that h=88.2353 and k=-186.176

OpenStudy (anonymous):

to solve for y, you first multiply through by 0.0085\[y-120 \approx 0.0085(x-88.2353)^2-(7785.467*.0085)\]then add 120\[y \approx 0.0085(x-88.2353)^2+53.8235\]now this is in the so-called standard form for a quadratic function from which we can read the vertex: V(88.2353, 53.8235)

OpenStudy (anonymous):

the standard form of a quadratic function with vertex (h, k) is\[y=a(x−h)^2+k\]compare this to the equation from your problem and you can see that h=88.2353 and k=53.8235